Fermi National Laboratory
 

Accelerator Update for May 9, 2007 - May 11, 2007

 

Wednesday May 9

The day shift began with the Tevatron (TeV) in a dry squeeze, with Operations monitoring the Recycler stash (241.1E10) and the Antiproton Source (Pbar) stack (11.3E10), with Pbar conducting reverse proton studies, with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson MTest experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations put the TeV into a shot setup at 9:07 AM.

LCW technicians reported at 9:37 AM that they were changing filters in the Meson MS2 area LCW system.

At 11:13 AM, Operations reported that beam wasn’t making it past Linac’s first RF Cavity station (LRF1). Experts investigated and at 11:34 AM reported that they had found a quadrupole (L:IQTM1U) power supply that needed its output raised by 50 amps.

Operations established store 5415 at 12:56 PM with an initial luminosity of 179.0E30.

The Neutron Therapy Facility (NTF) reported that they had treated their last patient for the day at 12:58 PM.

A MI expert began a MI-8 collimator study at 1 PM.

At 1:01 PM, LRF3 tripped off for the first of two times in about 85 minutes. It reset.

MTest resumed taking beam at 1:07 PM.

At 2:29 PM, BRF5 tripped off multiple times. Operators bypassed the station, contacted experts, and turned on BRF17 in its place.

At 3:16 PM, a TeV sector D17 separator sparked.

Operations reported at 3:25 PM that BRF5 was back on line and BRF17 had been turned off.

Operations reported at 4:37 PM that a Recycler expert replaced a correction element power supply (CE603).

Recycler experts reported at 5:33 PM that they had completed rigging in a replacement power supply for R:VDPA1 and that it had successfully passed a critical device test.

Pbar resumed stacking at 5:52 PM.

At 6:55 PM, Operations reported that they were unable to establish protons into the Recycler. They reported a problem with VDPA1. A power supply expert was called in.

At 8:08 PM, a power supply expert reported finding a LCW leak in VDPA1. The leak was on a braze joint and he said it would take a while to repair.

At 10:14 PM, Operations reported that a TeV sector B0/BA pond water pump had tripped off and caused cold compressor #4 at BA to trip off as well. The Duty Mechanic reported that there was an open in the pump’s winding. The problem would be addressed on Thursday.

Operations reported at 11:18 PM that the Recycler power supply VDPA1 was again operational, but not getting to its proper current. The power supply expert reported at 11:47 PM that the power supply was not tapped properly.

Thursday May 10

The midnight shift began with Operations monitoring store 5415 (58.21E30) and the stash (13.3E10), with Pbar stacking (112.2E10), and with NuMI and MiniBooNE taking beam.

Operations established a proton beam to the Recycler a little after midnight.

At 4:08 AM, store 5415 aborted at TeV sector F4. There was no quench. A QPM pulled a fast abort. Operations reported that the same kind of abort had happened back on 4/27/07 at A4.

Operations turned off the TeV at 4:46 AM and then prepared the TeV for access.

Operations prepared NuMI MI-65 for access at 5:15 AM.

Operations turned off Booster at 5:56 AM in preparation for an access at 6:20 AM.

Operations reported at 6:21 AM that R:VDPA1 had tripped off on an interlock sum fault. Operators contacted a power supply expert.

Operations turned off MI at 7:05 AM.

Interlock technicians reported at 7:20 AM that they had finished their safety system tests at B0.

CDF and D0 accessed their collision halls at 7:38 AM.

TeV experts began work on the sector A1 QPM problem at 8:04 AM.

Operations reported at 10:19 AM that the LCW pump at TeV sector B1 was replaced.

Interlock techs reported at 10:27 AM that they had completed their safety tests for MI.

NTF reported at 12:02 PM that they had treated their last patient for the day.

CUB techs began the heat exchanger switch at 1:11 PM for the Linac Klystron system. This took about 30 minutes. Then they changed the pumps for the MI chilled water system.

Operations reported at 2:35 PM that the TeV safety system tests had been completed.

Operations reported at 4:10 PM that the Booster safety system tests had been completed.

At 4:30 PM, cryo system techs reported a power glitch that tripped off the TeV sector A3 cold compressors. Operations reported that it also tripped off the MI RF system and a few other devices. The High Voltage Coordinator reported that ComEd admitted to a power glitch in their system.

Operations established beam to the Linac absorber at 5:20 PM.

At 5:30 PM, Operations reported the following problems: a failed Booster power supply (B:CORSUP), no readbacks for the MI-40 power supplies, and the TeV suffered a hard ground fault.

Operations prepared the TeV transfer hall and F sector for an access by 5:58 PM.

TeV experts reported at 6:32 PM that they found the ground fault at sector E43; a copper tubing line was crossing the conning tower safety leads.

An expert helped Operations get around the lack of readbacks at MI-40 and establish a MI ramp at 6:50 PM. He continued to investigate the readback problem.

Operations established beam to MiniBooNE at 7:13 PM.

Operations established beam to MI and NuMI at 7:41 PM.

A power supply expert reported at 7:44 PM that the Recycler’s VDPA1 was back online.

Operations established beam to Pbar at 7:53 PM.

Operations put the TeV through a dry squeeze at 9:41 PM.

Operations injected protons into the TeV at 10:43 PM for studies.

At 11:10 PM, the Tevatron electrical safety system tripped off. Operators called in an interlock tech. This put a halt to the TeV studies.

Friday May 11

The midnight shift began with an interlock tech investigating a TeV safety system problem, with Operations monitoring the stash (115.0E10), with Pbar stacking (10.4E10), and with NuMI and MiniBooNE taking beam.

The interlock tech reported at 1:28 AM that he had repaired a bad keytree switch in the MCR.

At 1:33 AM, BRF3 tripped off for the first of two times in 23 minutes. It reset.

Operations put the TeV through a dry squeeze at 3:28 AM.

Pbar began stacking around 3:50 AM.

Operations put the TeV through a wet squeeze at 4:08 AM.

At 4:35 AM, the beam fell out of the TeV. Operators investigated and discovered that the Tevatron Electron Lens (TEL2) had tripped off due to a vacuum interlock. The beam fell out of the TeV when TEL2 was turned back on.

Operations put the TeV into another wet squeeze at 5:29 AM.

At 5:43 AM, Operations reported an indication of a Pbar transport LCW leak. Operators called in experts.

Operations stashed antiprotons into the Recycler at 6:17 AM.

Operations prepared the Pbar transport area and rings for accesses at 6:38 AM.

Operations loaded a 14x0 wet squeeze into the TeV at 6:38 AM.

Operations loaded an energy conservation timeline into the Time Line Generator (TLG) at 6:50 AM.

Operations reported at 7:45 AM that the Pbar transport LCW leak was due to a bad hose on IQ7.

Operations established beam to NuMI and MiniBooNE at 7:49 AM.

Operations reported at 7:57 AM that the Recycler had stashed twice during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to establish a store this morning and then stack, stash, and store throughout the weekend.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/11/2007    email Fermilab